OpenStudy (anonymous):

Match each expression with its solution. Match Term Definition positive thirty−nine minus negative twenty−six A) negative thirteen negative thirty−nine minus positive twenty−six B) negative sixty−five positive thirty−nine minus positive twenty−six C) positive thirteen negative thirty−nine minus negative twenty−six D) positive sixty−five

OpenStudy (anonymous):

help please

OpenStudy (anonymous):

1) 39 - -26 = 39+26 = 65 = D 2)-39-26 = -65 = B 3)39 - 26 = 13 = C 4)-39 - -26 = -39 + 26 = -13 = A

OpenStudy (anonymous):

thanks you

OpenStudy (anonymous):

No problem